OpenStudy (anonymous):

Transform the equation into an equivalent equations in the standard form ax^2+bx+c=0 1)x^2+9x=10 2)x^2=3x-8

OpenStudy (anonymous):

I don't know how to begin

OpenStudy (anonymous):

are they asking you for the values or a b and c?

OpenStudy (anonymous):

if they do then rearrange the 1 and 2 equations into the first form and see which one it is

OpenStudy (anonymous):

Let me check my textbook

OpenStudy (anonymous):

1 and two are different questions

OpenStudy (anonymous):

their asking for no values

OpenStudy (anonymous):

Oh. 1) x^2+9x-10=0 so, then a=2, b=9, c=-10 2)x^2-3x+8=0 a=2, b=-3, c=8 I think this is what they mean? Without the values

OpenStudy (anonymous):

So, the top of the answers are the standard formally written answers

OpenStudy (anonymous):

I'm confused?

OpenStudy (anonymous):

do they want to make them into standard form?

OpenStudy (anonymous):

Yes, so that the equation = 0

OpenStudy (anonymous):

so, bring everything to one side, and leave 0 at the other

OpenStudy (anonymous):

Sounds like they just want you to mess around with the equations until it looks like ax^2+bx+c=0

OpenStudy (anonymous):

How do I do that I learn it today

OpenStudy (anonymous):

for 2)x^2=3x-8 x^2=3x-8 x^2-3x=-8 x^2-3x+8=0 You just bring everything over to the other side to make the eqation equal 0 like I just did.

OpenStudy (anonymous):

so how do we do ti for one?

OpenStudy (anonymous):

You try, and follow the steps that I gave you! :) Let's see if it works

OpenStudy (anonymous):

x^2+9x-10=0 add 10 to 0

OpenStudy (anonymous):